# Skip to the Good Part: Representation Structure & Inference‑Time Layer Skipping in Diffusion vs. Autoregressive LLMs

### 08 Mar 2026 v1

@@ Abstract
Autoregressive (AR) language models form representations incrementally through left‑to‑right prediction, whereas diffusion language models (dLLMs) are trained via full‑sequence denoising. Although recent dLLMs match AR performance, it remains unclear whether diffusion objectives fundamentally reshape internal representations across depth. We perform the first layer‑ and token‑wise representational analysis comparing native dLLMs (LLaDA), native AR models (Qwen2.5), and AR‑initialized dLLMs (Dream‑7B). We find that diffusion objectives result in different, more hierarchical abstraction with substantial early‑layer redundancy and reduced recency bias, while AR objectives produce tightly coupled, depth‑dependent representations. Critically, AR‑initialized dLLMs retain AR‑like representational dynamics despite diffusion training, revealing persistent initialization bias. Leveraging this observed representational redundancy, we introduce a static, task‑agnostic inference‑time layer‑skipping method requiring no architectural changes or KV‑cache sharing. Native dLLMs achieve up to 18.75% FLOPs reduction while preserving over 90% performance on reasoning and code generation benchmarks, whereas AR models degrade sharply under comparable skipping. These results link training objectives to representational structure and enable practical, cache‑orthogonal efficiency gains.
Machine Learning, ICML
1 Introduction
Figure 1: Representational redundancy enables efficient inference-time layer skipping. We evaluate zero-shot layer pruning across three architectures: LLaDA (diffusion LLM), Dream7B (dLLM initialized from Qwen2.5), and Qwen2.5-7B (AR-LLM). The diffusion-based LLaDA exhibits remarkable robustness, retaining 88.24% performance at 18.75% FLOPs reduction (6 layers skipped), demonstrating significant representational redundancy. Conversely, autoregressive models show brittle behavior with only 64.71% retention at 7.14% FLOPs reduction (2 layers), revealing concentrated, non-redundant representations. Top-right region indicates optimal performance-efficiency trade-off.
Transition from Next‑Token Prediction (NTP) to Diffusion: Autoregressive (AR) LLMs trained with NTP build predictions left‑to‑right, while diffusion LLMs (dLLMs) learn by iteratively denoising full‑sequence states. In text, this shift is often instantiated through discrete diffusion formulations that replace the AR factorization with stepwise refinement over the entire sequence. The knowledge gap: although recent dLLMs (e.g., LLaDA (Nie et al., 2025), DiffuCoder (Gong et al., 2025)) now match AR performance, we do not yet understand how their representations differ. Does looking at the full sequence during training actually change how features are abstracted and consolidated across depth/steps?
While the dominant focus, motivation in dLLM research has been on inference-time efficiency—parallel decoding, KV-caching, and architectural optimizations—the internal representational properties of these models remain largely unexplored. Recent work (Gong et al., 2025) has begun examining local versus global representations in dLLMs and demonstrated their potential for diverse exploration in RL-based fine-tuning contexts. However, a systematic analysis of how diffusion objectives shape representational abstraction across layers, and how this compares to AR models, is still missing. Understanding these representational dynamics is crucial, as they may reveal fundamental differences in how models organize information—analogous to how residual connections in ResNets enable different optimization landscapes compared to standard feedforward networks.
We hypothesize that training objective and curriculum materially shape an LLM’s internal geometry, leading diffusion‑trained models to organize information differently from AR models. In particular, the full‑sequence feedback in diffusion may encourage earlier or more global abstraction, with practical consequences for inference‑time redundancy. While architectural approaches like YOCO reduce memory/latency via cache‑once designs and early‑exit during prefill, our goal is orthogonal: can objective‑induced redundancy alone support inference‑only layer skipping—without any KV‑cache sharing or parameter tying?
To isolate the role of objective and initialization, we compare three families: a native dLLM (LLaDA), a native AR model (Qwen2.5), and an AR‑adopted dLLM (Dream‑7B) that initializes from Qwen2.5, providing a strong test of initialization bias. Our representational analysis across layers and tokens/denoising steps shows that Dream‑7B aligns more closely with Qwen2.5 than with LLaDA, indicating that AR initialization imprints persistent structure even after diffusion training. Motivated by this, we assess whether the observed redundancy can be directly leveraged via static, task‑agnostic layer skipping at inference time—without KV sharing—to yield measurable speedups with minimal performance loss, thereby complementing cache‑centric methods like YOCO. Below we summarize our contributions in this work:
@@ 1 Introduction
  • Representational analysis revealing objective-induced redundancy: We present the first systematic layer-wise and token-wise similarity analysis comparing native dLLMs, AR models, and AR-initialized dLLMs. We demonstrate that diffusion objectives induce stronger hierarchical abstraction with concentrated early-layer redundancy and minimal recency bias, while AR objectives maintain incremental token-by-token refinement throughout depth with strong recency bias. We further reveal strong initialization bias: AR-initialized dLLMs (Dream-7B) retain AR-like representational patterns despite diffusion training.
@@
  • Inference‑time layer skipping: Leveraging objective‑induced representational redundancy, we introduce a static, task‑agnostic layer‑skip policy that requires no KV‑cache sharing and no architectural modifications. Native dLLMs achieve up to 18.75% FLOPs reduction with minimal accuracy loss ($10\%$ degradation on average), while AR models show brittleness.
@@
  • Cross‑domain benchmarking We evaluate across reasoning (GSM8K, MATH-500) and code synthesis (HumanEval, MBPP) tasks, demonstrating consistent patterns: native dLLMs tolerate aggressive layer skipping (6+ layers with $90\%$ retention), AR-initialized dLLMs show intermediate robustness (2-4 layers), and native AR models are brittle (substantial degradation at 2 layers).
@@ 2 Layer-wise and Token-wise Similarity Analysis
Figure 2: Layer-skip mechanism for dLLMs. At each denoising step, high-similarity layers (shaded) are bypassed, with hidden states passed directly to the next active layer. This reduces per-step FLOPs while preserving the coarse-to-fine abstraction hierarchy.
@@
We hypothesize that this objective-level difference manifests as measurable representational redundancy that can be exploited for inference-time efficiency gains without architectural modifications or KV-cache sharing. Given that ‘coarse-to-fine’ generation is a common motivation and hypothesis for diffusion models, we take a first step in analyzing the rate of change of representations (within the shared embedding space) across both layers and tokens as illustrated in Fig. 5 and Fig. 6 for cosine similarity across tokens.
Methodology: We track the cosine similarity between consecutive layer representations $\mathbf{h}_{\ell}$ and $\mathbf{h}_{\ell+1}$ across all tokens in a sequence. For dLLMs, we compute this similarity at multiple denoising steps $t\in\{1,\ldots,T\}$; for AR models, we compute it during standard forward passes. We aggregate statistics across diverse prompts from our evaluation benchmarks.
1) Coarse-to-fine abstraction in native dLLMs: For LLaDA, from Fig. 5 the layer-wise similarity pattern remains largely consistent across denoising steps, suggesting a pronounced representational abstraction hierarchy. Early layers establish coarse representations with high inter-layer similarity (plateau regions with cosine similarity >0.95>0.95), while later layers and denoising steps perform iterative refinement rather than restructuring the representation at each step. This hierarchical organization indicates potential redundancy in early layers that can be exploited at inference time.
@@
Our representational similarity analysis complements the behavioral analysis in (Gong et al., 2025), which measures AR-ness through generation patterns (local consecutive next-token prediction and global earliest-mask selection). While their metrics capture output-level generation strategies—whether models follow left-to-right filling patterns—our layer-wise and token-wise cosine similarity analysis reveals internal representational dynamics: how hidden states evolve across depth and tokens.
@@
Recency Bias & Representational Abstraction Diffusion objectives reduce recency bias and promote hierarchical abstraction: LLaDA shows minimal recency bias with global representations across tokens, while AR models (Qwen2.5, Dream-7B) exhibit strong recency bias at all layers. This suggests diffusion training encourages coarse-to-fine abstraction, whereas AR training maintains incremental, token-by-token updates throughout network depth.
@@
As shown in Figure 3, Dream-7B’s average token-wise cosine similarity across layers follows Qwen2.5’s pattern remarkably closely throughout the network depth, despite undergoing diffusion training. In contrast, LLaDA exhibits a distinctly different profile: it begins with very high similarity (>0.95>0.95) in early layers, indicating redundant representations with smooth transitions, then transitions to lower similarity in later layers where refinement occurs. This hierarchical pattern—high redundancy in early layers, active refinement in later layers—is characteristic of native diffusion training and absent in AR-initialized models. This highlights the strong regularization effect of initialization on resulting representations and abstractions, persisting even after significant fine-tuning with diffusion-based objectives.
@@
Figure 3: Average token-wise cosine similarity across layers and denoising steps. LLaDA (native dLLM) exhibits high similarity (>0.9>0.9) in early layers with smooth transitions, followed by lower similarity in later layers where refinement occurs. Dream-7B closely follows Qwen2.5’s pattern despite diffusion training, revealing persistent initialization bias. Shaded regions show standard deviation across denoising steps for LLaDA and Dream-7B.
@@
Figure 5: Layer-wise cosine similarity across models all tokens decoded. Each row shows similarity between consecutive layers for (top) LLaDA, (middle) Qwen2.5, and (bottom) Dream-7B. High-similarity regions (yellow) indicate representational redundancy. Dream-7B’s pattern closely resembles Qwen2.5 despite diffusion training, revealing strong initialization bias.
@@
Magnitude Evolution One potential limitation of cosine similarity is its invariance to magnitude. To ensure our redundancy findings are not artifacts of magnitude collapse, we analyze the $\ell_{2}$ norm of hidden states across layers as observed in Fig. 7.We observed that the magnitude evolution is small for initial 60-70% layers and then rises steeply. There is also presence of sink tokens – super high magnitude than the rest of the tokens, as discussed in (Rulli et al., 2025).
@@ 3 Layer-skip at Inference
The observed high-similarity plateaus suggest that certain layers contribute minimally to representational transformation. We hypothesize that skipping these layers at inference time can reduce computational cost with minimal impact on task performance. We demonstrate with and without layer-skipping and prefix caching in Fig. 2

### 27 Apr 2026 v2

# A Comparative analysis of Layer-wise Representational Capacity in AR and Diffusion LLMs

@@ Abstract
Autoregressive (AR) language models build representations incrementally via left‑to‑right prediction, while diffusion language models (dLLMs) are trained through full‑sequence denoising. Although recent dLLMs match AR performance, whether diffusion objectives fundamentally reshape internal representations remains unclear. We perform the first layer‑ and token‑wise representational analysis comparing native dLLMs (LLaDA), native AR models (Qwen2.5), and AR‑initialized dLLMs (Dream‑7B), using cosine similarity across layers and tokens alongside static inference-time layer-skipping as an analytical probe of redundancy. We find that diffusion objectives produce more global representations with substantial early‑layer redundancy and reduced recency bias, while AR objectives yield tightly coupled, locally-structured representations. AR‑initialized dLLMs retain AR‑like dynamics despite diffusion training, revealing persistent initialization bias. Leveraging this redundancy, native dLLMs absorb up to 18.75% FLOPs reduction: retaining over 90% performance on math-reasoning and coding benchmarks, while AR models collapse under identical skipping, revealing that diffusion objectives, rather than architecture alone, induce depth redundancy that enables principled compression
From Next-Token Prediction to Diffusion Objectives. Autoregressive (AR) language models are trained via next-token prediction (NTP), constructing representations incrementally through left-to-right factorization. In contrast, diffusion language models (dLLMs) replace this causal factorization with full-sequence denoising, iteratively refining an entire token sequence from noise to data. Recent discrete diffusion models—such as LLaDA Nie et al. (2025) and DiffuCoder Gong et al. (2025)—have demonstrated performance on par with strong AR baselines across a range of downstream tasks. Despite this progress, a fundamental question remains unanswered: does training with a diffusion objective materially change how language models organize and abstract information internally, or are diffusion models representationally similar to AR models once performance is matched?
The representational gap. Most prior work on diffusion LLMs has focused on efficiency advantages—parallel decoding, verifier-based sampling, or architectural optimizations—rather than the structure of their learned representations. As a result, we lack a systematic understanding of how diffusion objectives shape internal geometry across depth and tokens, and how this differs from AR training. While recent studies Gong et al. (2025) have begun to explore local versus global behaviors in diffusion models, a comprehensive, layer- and token-wise comparison between AR and diffusion objectives is still missing. Such an analysis is crucial: representational structure reflects not only inference behavior but also latent inductive biases introduced by training objectives, much like how residual connections or normalization schemes influence optimization and generalization in deep networks.
Global vs. local representations. We hypothesize that the training objective itself—and in particular, whether supervision is delivered through causal next-token prediction or full-sequence denoising—fundamentally shapes representational abstraction in LLMs. We use the term global representations to denote hidden states that integrate information from the full input sequence, rather than being dominated by the most recent tokens. Full-sequence feedback in diffusion training naturally encourages such global integration, whereas AR training enforces a strong recency bias due to its causal structure. In particular, full-sequence denoising may encourage earlier consolidation of global abstractions, concentrating representational redundancy in shallower layers—consistent with recent findings that intermediate layers can encode richer representations than final layers Skean et al. (2025). To operationalize this distinction, we analyze representational similarity across layers and tokens, measuring how hidden states vary with token position and depth.
Isolating objective from initialization. A central challenge in comparing AR and diffusion models is disentangling the effects of training objective from architecture and initialization. To address this, we study three families: (i)a native AR LLM trained exclusively with NTP (Qwen2.5(Yang et al., 2024)), (ii)a native dLLM trained from scratch with a diffusion objective (LLaDANie et al. (2025)), and (iii)an AR-initialized dLLM that undergoes diffusion training after AR pretraining (Dream-7B(Ye et al., 2025b), initialized from Qwen2.5). This design allows us to isolate whether observed representational properties arise from the diffusion objective itself or are inherited from AR initialization.
Layer skipping as a diagnostic probe. Beyond static analysis, we introduce inference-time layer skipping as a controlled intervention to probe representational redundancy. Importantly, we do not propose layer skipping as a deployment strategy; rather, we use it as an analytical tool to stress-test how much computation can be removed before representations and performance degrade. If early layers are redundant or globally integrated, their removal should incur limited performance loss; conversely, models with tightly coupled, locally structured representations should fail under the same perturbation. This static, task-agnostic policy requires no KV-cache sharing and no architectural modifications, providing an orthogonal complement to cache-centric methods such as YOCO Sun et al. (2024).
Our findings. Across all analyses, a consistent picture emerges. Native dLLMs learn globally coherent representations with substantial early-layer redundancy, exhibiting reduced recency bias and high cross-token similarity in early layers. AR-initialized dLLMs, by contrast, retain AR-like representational dynamics even after diffusion training, demonstrating a persistent imprint of initialization and confirming that global redundancy is not a trivial consequence of architecture alone. These representational differences translate directly into measurable efficiency gaps: native dLLMs tolerate up to 18.75% FLOPs reduction via layer skipping while retaining over 90% performance, whereas AR models collapse under identical interventions.
Contributions. We summarize our contributions as follows:
@@
  • Representational analysis revealing objective-induced redundancy. We present the first systematic layer-wise and token-wise similarity analysis comparing native dLLMs, AR models, and AR-initialized dLLMs. We show that diffusion objectives produce more global representations with concentrated early-layer redundancy and minimal recency bias, while AR objectives maintain incremental, locally structured refinement with strong recency bias throughout depth. We further reveal a strong initialization bias: AR-initialized dLLMs (Dream-7B) retain AR-like representational patterns despite diffusion training, aligning more closely with Qwen2.5 than with LLaDA.
@@
  • Inference-time layer skipping. Leveraging objective-induced representational redundancy, we introduce a static, task-agnostic layer-skip policy requiring no KV-cache sharing and no architectural modifications. Native dLLMs (LLaDA) achieve up to 18.75% FLOPs reduction with <<10% average accuracy degradation, while AR models show substantial brittleness under the same intervention, providing an architecture-agnostic complement to cache-centric designs.
@@
  • Cross-domain benchmarking. We evaluate across math reasoning (GSM8K, MATH-500) and code synthesis (HumanEval, MBPP), demonstrating consistent patterns: native dLLMs tolerate aggressive layer skipping (6 layers, >>90% average retention), AR-initialized dLLMs exhibit intermediate robustness ($\sim$78% average retention at 2-layer skip, declining at higher compression), and native AR models degrade substantially at 2-layer skip ($\sim$63% average retention on tasks with non-trivial baselines).
@@ 1 Layer-wise and Token-wise Similarity Analysis
We hypothesize that this objective-level difference manifests as measurable representational redundancy exploitable for inference-time efficiency gains without architectural modifications or KV-cache sharing. Specifically, we analyze the rate of change of representations across both layers and tokens—examining whether diffusion training tends to produce more global representations compared to the heavily local representations of AR models—as illustrated in Fig. 2 and Fig. 3.
Methodology: We track the cosine similarity between consecutive layer representations $\mathbf{h}_{\ell}$ and $\mathbf{h}_{\ell+1}$ across all tokens in a sequence. Formally, for token $i$ at layer $\ell$, we compute:
$\text{sim}(\mathbf{h}_{\ell}^{(i)},\mathbf{h}_{\ell+1}^{(i)})=\frac{\mathbf{h}_{\ell}^{(i)}\cdot\mathbf{h}_{\ell+1}^{(i)}}{\|\mathbf{h}_{\ell}^{(i)}\|\,\|\mathbf{h}_{\ell+1}^{(i)}\|}$ (1)
and aggregate across tokens and prompts to obtain a layer-wise similarity profile.
Why cosine similarity? We choose cosine similarity for its robustness to magnitude changes in hidden states. Because hidden-state norms can vary substantially across layers (as we show in Figure 4), a magnitude-sensitive metric would conflate representational change with scale drift. Cosine similarity isolates directional change in representation space, making it a more reliable indicator of whether a layer is performing meaningful transformation. This choice is further supported by recent work Men et al. (2025), which demonstrates that their proposed BI metric—operating on the same mathematical rationale as cosine similarity—is the most robust among four evaluated metrics for layer-wise representational analysis. While our methodology was developed independently, this strongly reinforces our choice.
We also considered metrics like Linear Centered Kernel Alignment (CKA), a natural alternative for measuring representational similarity, but it is unsuitable for our token-wise, sample-wise analysis. CKA is designed to operate at the dataset level: it relies on mean-centering across a batch of examples, meaning a true per-token kernel cannot be defined. Also Jiang et al. (2024) empirically demonstrates that cosine similarity closely tracks CKA in layer-wise analysis. We therefore use cosine similarity directly, which is both more interpretable and computationally efficient in our per-token setting.
For dLLMs, we compute this similarity at multiple denoising steps $t\in\{1,\ldots,T\}$; for AR models, we compute it during standard forward passes. We aggregate statistics across diverse prompts from our evaluation benchmarks.
1) Global representations and early-layer redundancy in native dLLMs: For LLaDA, from Fig. 2 the layer-wise similarity pattern remains largely consistent across denoising steps, suggesting a tendency toward more global representational abstraction. Early layers exhibit high inter-layer similarity (plateau regions with cosine similarity >0.95>0.95), while later layers and denoising steps perform iterative refinement. This organization—high redundancy in early layers, active refinement in later layers—indicates potential redundancy that can be exploited at inference time.
@@
Our representational similarity analysis complements the behavioral analysis in Gong et al. (2025), which measures AR-ness through generation patterns (local consecutive next-token prediction and global earliest-mask selection). While their metrics capture output-level generation strategies—whether models follow left-to-right filling patterns—our layer-wise and token-wise cosine similarity analysis reveals internal representational dynamics: how hidden states evolve across depth and tokens.
@@
Recency Bias & Representational Abstraction Diffusion objectives reduce recency bias and promote global representations: LLaDA shows minimal recency bias with global representations across tokens, while AR models (Qwen2.5, Dream-7B) exhibit strong recency bias at all layers. This suggests diffusion training encourages more global abstraction, whereas AR training maintains incremental, token-by-token updates throughout network depth Figure 1, 3
@@
As shown in Figure LABEL:fig:avg_token_similarity, Dream-7B’s average token-wise cosine similarity across layers follows Qwen2.5’s pattern remarkably closely throughout the network depth, despite undergoing diffusion training. In contrast, LLaDA exhibits a distinctly different profile: it begins with very high similarity (>0.95>0.95) in early layers, indicating redundant representations with smooth transitions, then transitions to lower similarity in later layers where refinement occurs. This hierarchical pattern—high redundancy in early layers, active refinement in later layers—is characteristic of native diffusion training and absent in AR-initialized models. This highlights the strong regularization effect of initialization on resulting representations and abstractions, persisting even after significant fine-tuning with diffusion-based objectives.
@@
Figure 2: Layer-wise cosine similarity across models 128 tokens decoded. Each row shows similarity between consecutive layers for (top) LLaDA, (middle) Qwen2.5, and (bottom) Dream-7B. High-similarity regions (yellow) indicate representational redundancy. Dream-7B’s pattern closely resembles Qwen2.5 despite diffusion training, revealing strong initialization bias.
@@
Magnitude Evolution One potential limitation of cosine similarity is its invariance to magnitude. To ensure our redundancy findings are not artifacts of magnitude collapse, we analyze the $\ell_{2}$ norm of hidden states across layers as observed in Fig. 4.We observed that the magnitude evolution is small for initial 60-70% layers and then rises steeply. There is also presence of sink tokens, super high magnitude than the rest of the tokens, as discussed in Rulli et al. (2025).
@@ 2 Layer-skip at Inference
The observed high-similarity plateaus suggest that certain layers contribute minimally to representational transformation. We hypothesize that skipping these layers at inference time can reduce computational cost with minimal impact on task performance. We demonstrate with and without layer-skipping and prefix caching in Fig. LABEL:fig:dllm-layer-skip
@@
Algorithm 1 Static Layer-Skip Selection Psuedo Code
1:Hidden states $\{\mathbf{h}_{\ell}\}_{\ell=0}^{L}$, number of layers to skips $n_{\text{skip}}$
2:Skip set $\mathcal{L}_{\text{skip}}$
3:Compute $s_{\ell}=\cos(\mathbf{h}_{\ell-1},\mathbf{h}_{\ell})$ for $\ell=1,\ldots,L$ $\triangleright$ layer-wise similarity
4:Sort layers by $s_{\ell}$ in descending order
[…]
6:for each layer $\ell$ in sorted order do
7:  if and $\ell\pm 1\notin\mathcal{L}_{\text{skip}}$ and $|\mathcal{L}_{\text{skip}}|<n_{\text{skip}}$ then
8:   $\mathcal{L}_{\text{skip}}\leftarrow\mathcal{L}_{\text{skip}}\cup\{\ell\}$
9:  end if
10:end for
11:return $\mathcal{L}_{\text{skip}}$ $\triangleright$ see Appendix for full implementation details
3 Experimental Setup
Models.
We evaluate three families to disentangle training objective and initialization: (i) a native diffusion LLM, LLaDA (we use the 8B Base / Instruct checkpoints) Nie et al. (2025); (ii) a native autoregressive (AR) model, Qwen2.5 (7B Base / Instruct) Yang et al. (2024); and (iii) an AR-initialized diffusion LLM, Dream-7B (Instruct) (Ye et al., 2025a). Unless stated otherwise, all models are evaluated with their public inference code and default tokenizers.
Benchmarks.
Prompting and answer extraction.
For GSM8K, we use a few-shot rationale prompt with an explicit Final Answer: line; we strip formatting and compare normalized numbers. For HumanEval and MBPP, we request a single Python function and evaluate with the official test suites; we report pass@1 and pass@k. The same prompts are used across models to ensure comparability. We follow the exact inference setting of Chen & Liu (2026), including the use of prefix caching to mirror the Fast-dLLM evaluation setting, ensuring accurate and comparable evaluations across models.
Decoding & sampling.
[…]
Layer-skipping evaluation (ours).
4 Results
Native dLLMs Enable Aggressive Layer Skipping. Table 1 and Figure LABEL:fig:combined show that native diffusion LLMs are markedly more robust to layer skipping than autoregressive models. For LLaDA, skipping 6 layers (18.75% FLOPs reduction) preserves 93–97% of baseline performance across GSM8K, MATH500, and MBPP, with HumanEval slightly lower at 83%. Even at an 8-layer skip (25% FLOPs reduction), performance retention remains high (60–94%) as shown in Appendix Table 3, placing LLaDA firmly in the favorable efficiency–quality regime (top right of Figure LABEL:fig:layer-skip-flops).
In contrast, autoregressive models degrade rapidly and unevenly under similar interventions. Qwen2.5 shows highly task-dependent sensitivity: skipping only 2 layers (7.14% FLOPs reduction) reduces GSM8K retention to 48% and MBPP to 62%. We note that retention percentages should be interpreted alongside the absolute accuracies shown in parentheses: apparent gains such as Qwen2.5’s 143% retention on MATH500 at a 2-layer skip are not meaningful, as the baseline absolute accuracy is only 0.07—an increase to 0.10 represents 3 additional correct answers out of 100 samples, well within sampling noise. Dream-7B-base, when layer skipping is applied with optimized non-consecutive layer selection, exhibits noticeably stronger robustness than Qwen2.5 at low skip rates—matching LLaDA on GSM8K (101% vs. 95%) and HumanEval (95% vs. 96%) at a 2-layer skip. However, Dream-7B-base diverges from LLaDA at higher skip counts: at 4 layers skipped, retention drops to 50–80% (vs. 87–100% for LLaDA), and falls to 8–25% at 6 layers. This suggests that while diffusion fine-tuning does partially restructure representational redundancy at low compression rates, the underlying representations remain more brittle than those of natively trained dLLMs under aggressive layer pruning.
[…]
Table 1: Performance comparison across models and layer skipping configurations. Values represent retention percentages relative to the 0-layer baseline, with absolute accuracy shown in parentheses for baseline rows. Retention values should be interpreted alongside absolute accuracies: large percentage swings on tasks where baseline accuracy is near floor (e.g., $\leq$0.10) reflect sampling noise rather than meaningful robustness, as small absolute fluctuations produce arbitrarily large percentage changes.
Task Layers Skipped Models
LLada-8B-Instruct Dream-7B-base Dream-7B-Instruct Qwen2.5-7B
GSM8K 0 (baseline) 100% (0.79) 100% (0.74) 100% (0.80) 100% (0.21)
2 95% (0.75) 101% (0.75) 91% (0.73) 48% (0.10)
4 94% (0.74) 68% (0.50) 18% (0.14) 24% (0.05)
6 94% (0.74) 22% (0.16) 4% (0.03) -
MATH500 0 (baseline) 100% (0.38) 100% (0.36) 100% (0.48) 100% (0.07)
2 108% (0.41) 69% (0.25) 77% (0.37) 143% (0.10)
4 87% (0.33) 50% (0.18) 31% (0.15) 100% (0.07)
6 97% (0.37) 25% (0.09) 0% (0.00) 0% (0.00)
HumanEval 0 (baseline) 100% (0.53) 100% (0.64) 100% (0.62) 100% (0.46)
2 96% (0.51) 95% (0.61) 90% (0.56) 78% (0.36)
4 91% (0.48) 61% (0.39) 29% (0.18) 33% (0.15)
6 83% (0.44) 8% (0.05) 0% (0.00) 7% (0.03)
MBPP 0 (baseline) 100% (0.56) 100% (0.44) 100% (0.75) 100% (0.63)
2 107% (0.60) 48% (0.21) 80% (0.60) 62% (0.39)
4 100% (0.56) 80% (0.35) 37% (0.28) 27% (0.17)
6 93% (0.52) 14% (0.06) 12% (0.09) 19% (0.12)
4.1 Layer Distribution and Skip Sensitivity
Tables 2 reveal that consecutive layer skipping is catastrophic. For LLaDA at 8-layer skip, allowing consecutive removal drops GSM8K retention from 94% to 40% and HumanEval from 60% to 8%. Similarly, accuracy drops a lot when skipping 6-8 layers uniformly or randomly justifying the utility of our layer-skip algorithm.
Our algorithm avoids this by maintaining representational continuity. Analysis and Fig.5 shows skipped layers concentrate in early network depth (first 40–60%), aligning with our observation that early layers tend to develop more global representations with high redundancy, while later layers perform critical fine-grained refinement.
Table 2: Comparison of Layer Skipping Strategies (Accuracy Retention) for LLada-8B-Instruct where Top-K is consecutive layers, Unif skips layers uniformly, and Rand skips random layers
Skip GSM8K HumanEval
Ours Top-k Unif Rand Ours Top-k Unif Rand
6 94% 95% 87% 77% 83% 70% 55% 36%
8 94% 40% 71% 62% 60% 8% 30% 21%
Figure 5: Which layers are skipped? (6-layer skip) Distribution over layer indices on LLaDA (HumanEval). Disallowing consecutive skips concentrates selections in early layers (first 40–60%).
5 Related Work
Diffusion Language Models and Representations. Diffusion language models (dLLMs) replace autoregressive decoding with bidirectional denoising objectives, enabling parallel decoding and global context modeling. Foundational work on discrete diffusion Austin et al. (2021) led to recent dLLMs such as SEDD Lou et al. (2023), and LLaDA Nie et al. (2025); Bie et al. (2025), which achieve competitive language modeling performance. Dream‑7B Ye et al. (2025a) adapts pretrained AR models to diffusion training, while MDLM Sahoo et al. (2024) simplifies diffusion objectives. Despite this progress, how diffusion objectives shape internal representations—especially relative to AR and AR‑initialized models—remains insufficiently understood.
Additionally, efforts have been made to integrate KV caching mechanisms to reduce redundant computation Ma et al. (2025); Liu et al. (2025). An alternate line of work focuses on step distillation of dLLMs Deschenaux & Gulcehre (2025); Qian et al. (2026) towards accelerating dLLM inference.
Representational Structure and Initialization Effects. Prior work shows that language models organize representations hierarchically across depth, with earlier layers capturing coarse features and deeper layers refining task‑specific abstractions Jawahar et al. (2019). While representational dynamics in AR models have been studied, systematic analyses comparing AR and diffusion models are rare. Our work provides direct, layer‑ and token‑level evidence of this initialization bias in AR‑adapted dLLMs and contrasts it with the representational redundancy that emerges in native diffusion models.
Appendix A Appendix
Appendix B Summary and Future Work
We presented the first systematic layer- and token-wise representational analysis comparing native dLLMs, AR models, and AR-initialized dLLMs. We find that diffusion objectives produce more global representations with substantial early-layer redundancy and reduced recency bias, while AR objectives yield tightly coupled, locally-structured representations, and AR initialization imprints this structure persistently even after diffusion training.
Deeper representational analysis: Cosine similarity is a first-order probe, it captures directional stability across layers but does not characterize the intrinsic information content of representations.
Broader model coverage and training objectives. Our analysis covers one native dLLM (LLaDA) and one AR-initialized dLLM (Dream-7B), leaving open whether the global representation pattern generalizes across the diffusion model family. Key open questions include: Does it extend to models with fundamentally different transition kernels, such as uniform diffusion (SEDD)? Does RL-based post-training—which has been shown to reshape representational geometry in AR models—further shift dLLMs toward or away from AR-like structure? Extending this analysis to post-trained dLLMs and multi-modal diffusion architectures would substantially broaden the scope and impact of these findings.
Appendix C Layer-skip Algorithm